The depth of the crater of a Volcano was calculated in the following manner. From a helicopter flying vertically upwards at 20 feet per second, a small bomb was released at the instant the helicopter was 120 feet above the crater surface. The sound of the explosion was heard 9 seconds later. The speed of the sound is 1000 fps. What is the depth of the crater in feet?



1
Expert's answer
2022-02-28T10:06:26-0500

"h=\frac{v_0+\frac{2gh}{v+gt_0}}t=250~ft."